OpenStudy (anonymous):

what is the zero of the function f(x)=3x-6

OpenStudy (anonymous):

will give medal

OpenStudy (anonymous):

i think a zero means where f(x) which is the output is equal to 0, try 0 = 3x - 6 and solve for x

OpenStudy (anonymous):

y = 3x - 6 3x = y + 6 x = (1/3) y + 2 f (x) = y g (y) = x g (y) = (1/3) y + 2 g (x) = (1/3) x + 2